# Numerical Aspects of Large Deviations

### v1

@@ Abstract
An introduction to numerical large-deviation sampling is provided. First, direct biasing with a known distribution is explained. As simple example, the Bernoulli experiment is used throughout the text. Next, Markov chain Monte Carlo (MCMC) simulations are introduced. In particular, the Metropolis-Hastings algorithm is explained. As first implementation of MCMC, sampling of the plain Bernoulli model is shown. Next, an exponential bias is used for the same model, which allows one to obtain the tails of the distribution of a measurable quantity. This approach is generalized to MCMC simulations, where the states are vectors of $U(0,1)$ random entries. This allows one to use the exponential or any other bias to access the large-deviation properties of rather arbitrary random processes. Finally, some recent research applications to study more complex models are discussed.
@@ 1 Motivation
As toy example for the following content, we use a very simple model, i.e., the Bernoulli experiment. This is a series of $n$ independent coin flips with outcomes $y_{i}\in\{0,1\}$, $\underline{y}=(y_{1},\ldots,y_{n})$. Let $\alpha$ be the probability of obtaining the result $1$ in a single coin flip, i.e.
@@ 2 Simple and Biased Sampling
The most natural implementation of a stochastic system is to sample the states or configuration according to the original distribution of the system. Performing measurements means to take just the configurations, calculate the measurable quantities, and then obtain averages or histograms. This is introduced in the first subsection of this chapter. Next, as an example the Bernoulli experiment is considered. Then it is shown that one can reach the rare events, i.e., the tails of the distributions, by using biased sampling. Finally, this will be illustrated for the Bernoulli experiment again.
@@ 2.2 Simple Sampling for the Bernoulli Process
The result obtained within a simple simulation, here for $n=50$ and $\alpha=0.3$, could look like shown in Fig. 1. We consider three different values for the number $M$ of samples, namely $M=10^{4}$, $10^{6}$ and $10^{6}$. The histograms follow the true distribution Eq. (3) better for increasing $M$, but the tails are not accessible.
Figure 1: Numerically measured histograms of the number $l$ of 1’s obtained from $n=50$ coin flips. The histograms are measured from $M=10^{4}$, $10^{6}$ and $10^{6}$ experiments. The solid line shows the analytical solution Eq. (3).
@@ 2.4 Biased Sampling for the Bernoulli Process
In order the address these two problems, we first have to be able to sample according to more or less arbitrary distributions with possibly unknown normalization factor. This can be achieved by using Markov chains, which are covered next.
@@ 3 Markov Chain Monte Carlo Simulations
Often, the most simple numerical estimation procedure is uniform sampling. Here one generates a certain number $M$ states $\{\underline{y}^{i}\}$ ($i=1,\ldots,M$) randomly, with uniform probability. A typical value of $M$ could be, for example, $10^{6}$. For the Bernoulli example, one would obtain the uniform sampling by simply choosing for all entries $y_{i}=0,1$ with probability 0.5, instead of the actual probabilities $(1-\alpha)$ and $\alpha$, respectively. To estimate the average $\langle A\rangle$, one calculates the observable for all sampled states and weights the results with the correct probabilities. One has to normalize this average, resulting in
$\langle A\rangle\approx\overline{A}^{(1)}:=\sum_{\underline{y}^{i}}A(\underline{y}^{i})P(\underline{y}^{i})/\sum_{\underline{y}^{i}}P(\underline{y}^{i})$ (12)
It can be easily checked that this way of normalizing is correct. Say the measurable quantity was boringly constant $A\equiv 1$, then Eq. (12) would result in $\sum_{\underline{y}^{i}}P(\underline{y}^{i})/$ $\sum_{\underline{y}^{i}}P(\underline{y}^{i})=1$ which is correct. Also, if the sample consisted of all $K$ possible states, then the sum $\sum_{\underline{y}^{i}}P(\underline{y}^{i})$ in the denominator is one by normalization and Eq. (12) reduces to the exact expectation value Eq. (11). Note also that for Eq. (12) it is sufficient to know the probabilities without normalization, since this cancels from the ratio.
@@
Note that $P$ denoted so far the original probabilities of the system. But, in the present context, it can be the biased probabilities $Q$ to drive the system to desired but unlikely configurations, as done for large-deviation sampling in section 2.3. In this case, one would have to chose $A$ appropriately to unbias the result, as discussed previously. We will come back to large-deviation sampling later on in section 4, but for the moment we assume just arbitrary given probabilities $P$ according which we desire to sample, let them be the original or biased ones.
@@ 3.1 Markov Chains
Under specific conditions [4], in particular if there is only one largest left eigenvalue $\lambda=1$ of the matrix $W$ the probability distribution will converge to a stationary, i.e. time-independent distribution
@@
This limit has to be independent of any given initial state $\underline{y}(0)$. The Markov process is called ergodic in this case. This basically means that one can reach each state from any other state by a finite sequence of transitions.
Note that there are non-ergodic systems, a simple example is shown in Fig. 4.
@@
In general and even if one takes the Metropolis-Hastings algorithm, there is a lot of freedom in choosing the transition probabilities, influencing strongly the efficiency of the corresponding algorithm. One general rule is, the more the configurations are changed between two time steps, the faster the algorithm will walk through configuration space, thus the more efficient the algorithm is. Unfortunately, these sytsem-wide changes are hard to set up, often one has to be satisfied with single-variable changes.
@@ 3.2 MCMC for the Bernoulli Process
As mentioned before, there are in general many possibilities to set up the transition probabilities in the matrix $W$. Here we proceed as follows: Given the current state $\underline{y}=\underline{y}(t)$, we just select $n_{\rm s}$ times a randomly chosen entry and redraw it according to the original probabilities. Therefore, we perform $n_{\rm c}$ new coin tosses, while the outcomes of the other coin tosses contained in $\underline{y}$ are not touched. This is implemented by the following algorithm:
@@
The resulting configuration is $\underline{y}(t+1)$. Each time a new index $i$ is drawn randomly and independently, this might result in repeated redrawing some entries. Since each coin toss is independent, only the last one will be effective for a given entry. Thus, the actual number of entries actually redrawn may be smaller than $n_{\rm s}$. A implmentation in C is contained in mc_bernoulli.c [^4] The code can be compiled in a Unix shell by cc -o mc_bernoulli mc_bernoulli.c -lm . in the function bernoulli_mc_step0().
@@
For obtaining the transition probabilities, we consider two states $\underline{y}$ and $\underline{z}$ that are connected by a step in the Markov chain. Let $I_{\rm c}$ contain those indices where $\underline{y}$ and $\underline{z}$ differ, i.e., $I_{\rm c}=\{i|y_{i}\neq z_{i};i=1,\ldots,n\}$. Note that $\underline{y}$ and $\underline{z}$ can have a non-zero transition probability only when $n_{\rm s}\geq|I_{\rm c}|$
Furthermore, let $I_{\rm s}$ be the actual set of indices selected by the algorithm, with $|I_{\rm s}|\leq n_{\rm s}$. This set will most of the time be a super set of $I_{\rm c}$, because for some of the selected entries redoing the coin flip might lead to the previous outcome. Hence, for the entries in $I_{\rm s}\setminus I_{\rm c}$, the entries are not changed, like those entries which do not appear in $I_{\rm s}$ anyway. We denote by $R(I_{\rm s})$ the probability that the algorithm selects a certain subset $I_{\rm s}$, which might be a complicated function of $I_{\rm s}$. For the entries $i$ which appear in $I_{\rm s}$, the joint probability for obtaining the outcomes $z_{i}$ is simply $\prod_{i\in I_{\rm s}}\alpha^{z_{i}}(1-\alpha)^{1-z_{i}}$. Splitting $I_{\rm s}$ into $I_{\rm c}$ and $I_{\rm s}\setminus I_{\rm c}$, and summing over all possible sets $I_{\rm s}$, this leads to the following transition probability:
@@
Figure 6: Markov chain for the Bernoulli case with $n=50$ coin flips and $\alpha=0.5$. Shown is the number $l(t)=\sum_{i}y_{i}(t)$ of 1’s, averaged over 1000 runs, starting at $l(0)=0$. Three different numbers $n_{\rm s}=1,2,5,10$ of entry changes are considered. The horizontal line indicates the expectation value $\alpha n$.
In Fig. 6 $l(t)$ is shown for different number $n_{\rm s}$ of selected entries for changes. The more entries are changed, the faster the convergence is. This is in general true for all MCMC algorithms: the approach is the more efficient, the faster it can move through configuration space.
@@ 3.3 Metropolis-Hastings algorithm
Note that for the opposite move $\underline{z}\to\underline{y}$ the vectors $\underline{z}$ and $\underline{y}$ exchange and therefore $Q$ is replaced by $1/Q$. This also means that for one of the two moves between $\underline{z}$ and $\underline{y}$ the transition probability is one. Hence, the Metropolis-Hastings choice is maximal!
@@ 3.4 Metropolis-Hastings Algorithm for the Bernoulli Process
Hence, one can calculate the acceptance probability without performing the actual flip beforehand. Thus, one performs a single-variable flip with probability $\tilde{W}$ to yield $\underline{y}(t+1)$, otherwise the current configuration $\underline{y}(t)$ is kept. A corresponding C Code is contained in mc_bernoulli.c in the function bernoulli_metropolis().
@@ 4 Biased Markov Chain Sampling
Now we come back to the main target, to obtain for a random process the distribution $Q(S)$ of a measurable quantity $S$, which we also call score. The distribution can be obtained in principle by considering all possible configurations $\underline{y}$ and adding up the probabilities $P({\underline{y}})$ of those configurations for which $S(\underline{y})$ has the desired value $S$, i.e.
@@ 4.1 Exponential Bias
where ${S}(\underline{y})$ is the score evaluated for the configuration. This bias appears very natural for physicists since it corresponds to the Boltzmann weight of the canonical ensemble. Therefore $\Theta$ is a temperature-like parameter, short just temperature, which controls the impact of the bias. Note that also other biases are possible, in principle everything that works is good. Nevertheless, for many systems the exponential bias works very well and it is also used for analytical calculations, where one can take advantage of biases to obtain tails of distributions in a similar way[7, 8].
@@
  • In contrast to real physical systems, also temperatures $\Theta<0$ make sense. Now, the bias is larger for configurations where $S=S(\underline{y})$ is larger than typical values of $S$.
@@
During the generation of configurations $\underline{y}$, one measures always the corresponding values $S=S(\underline{y})$, which finally can be collected in histograms which approximate the distributions $Q_{\Theta}(S)$ of $S$ in the biased ensemble at temperature $\Theta$. The results will depend of $\Theta$. This means that the histograms will cover a certain range of values of $S$. Therefore, one usually has to generate configurations for several values of $\Theta$ in order to obtain a full or at least a large range of values of $S$. Details how the final estimate for the distribution $Q(S)$ can be obtained from combining the result generated at different values of $\Theta$ is presented in Sec. 4.3.
@@ 4.2 Bias for the Bernoulli Process
For the Bernoulli process of $n$ coin flips, the MCMC algorithm redraws in each step a number $n_{\rm s}$ of randomly chosen elements of the current configuration $\underline{y}(t)$, with the original coin probability $\alpha$ as of Eq. (1). For the changed elements, the current values are stored in a vector, i.e. an array old[]. This allows to restore the previous configuration, if the trial configuration is rejected. Note that for restoring the configuration, the current values are restored in reverse order, to prevent a mistake if some elements are by chance selected more than once for a change.
@@
algorithm biased-MC-Bernoulli($\underline{y}$, $n$, $n_{\rm mc}$, $n_{\rm s}$, $\Theta$, function $S(\cdot)$)
@@
Note that in case of acceptance, nothing has to be done, since the trial state is stored directly in $\underline{y}$. The C code can be found in mc_bernoulli.c as function bernoulli_mc_step_bias().
@@
With the mc_bernoulli.c code we can generate data for $S_{3+}$, The Metropolis algorithm is chosen by the -alg 2 option. First, we simulate the almost unbiased case for which we choose $\Theta=1000$. For this case, a large number $n_{\rm s}=45$ of changes are performed in each step. This is stated with the -change option when calling the program. Here, we consider $n=51$ coin flips for $\alpha=0.5$, which are the third and second last arguments, respectively. The number of Monte Carlo steps is the last argument, 100000 here. The first 10000 steps are for equilibration and therefore ignored for measuring the histogram, as stated by the argument for the -histo option. When calling the program in Unix from a command line, this can be achieved by entering
@@
You should perform simulations yourself using the code and observe how varying the parameters, like the number $n_{\rm s}$ of redrawn coin flips, lead to changes of the results. In particular the empirical acceptance probability is affected, which is also written into the output file. It is also interesting to change $\Theta$ and try to sample the largest possible values, which is here $S_{3+}=13$.

### v2

@@ Abstract
An introduction to numerical large-deviation sampling is provided. First, direct biasing with a known distribution is explained. As simple example, the Bernoulli process is used throughout the text. Next, Markov chain Monte Carlo (MCMC) simulations are introduced. In particular, the Metropolis-Hastings algorithm is explained. As first implementation of MCMC, sampling of the plain Bernoulli model is shown. Next, an exponential bias is used for the same model, which allows one to obtain the tails of the distribution of a measurable quantity. This approach is generalized to MCMC simulations, where the states are vectors of $U(0,1)$ random entries. This allows one to use the exponential or any other bias to access the large-deviation properties of rather arbitrary random processes. Finally, some recent research applications to study more complex models are discussed.
@@ 1 Motivation
As toy example for the following content, we use a very simple model, i.e., the Bernoulli process. This is a series of $n$ independent coin flips with outcomes $y_{i}\in\{0,1\}$, $\underline{y}=(y_{1},\ldots,y_{n})$. Let $\alpha$ be the probability of obtaining the result $1$ in a single coin flip, i.e.
@@
The present text tries to be very comprehensive, including also very fundamental material, which many readers might know partially. Thus, a short overview is given now, to allow the reader maybe to skip some sections. In section two, simple sampling is introduced, which means that a probability distribution is sampled according its original underlying distribution. In the best case, this is done by direct sampling, where each call in the program to a corresponding function yields an independent realization of the stochastic process. This is applied to the Bernoulli process. Then biased sampling is introduced, which means that the system is sampled according to a distribution different from the original one. First, the educated variant is discussed, where one knows how one has to “push” a system by changing system parameters to sample it in the region of interest. Again the Bernoulli model is considered in test simulations, which shows than indeed the tails of the distributions can be reached.
In the third section, Markov chain Monte Carlo simulations are presented. First, Markov chains are introduced and the fundamental Master equation is explained. Next, a simple Markov chain for the Bernoulli process is built. Then, the Metropolis-Hasting algorithm is explained, which allows for sampling according to rather general desired distributions, which is in particular useful for cases where direct sampling does not work. Here again, the Bernoulli process is used as example, for pedagogical reasons, considering that direct sampling is possible there.
In the fourth section, the two main concepts, biasing and Markov chains, are combined. This allows for “blind” sampling, where the process under investigation is driven into the rare-event region of interest, without the need to know how exactly the system has to be “pushed”. This is done automatically by using a bias where the bias depends on the quantity of interest. Here, the standard case of an exponential bias, also called exponential tilt, is described. To demonstrate this, the Bernoulli process is considered. Here, unusual quantities of measurements, like the number of consecutive equal coin flips of length of at least 3 are used. This is useful, because the educated sampling does not work for such quantities. Then it is explained, how to obtain the true distribution, i.e., how to remove the bias from the sampled data. This involves also the determination of normalization constants, i.e., partition functions. In the last part, it is explained how the basic algorithm presented so far for the Bernoulli process can be generalized. This yields a “black box” approach, which means that it is applicable for rather arbitrary stochastic processes, at least if the quantity of interest is a scalar variable. Finally, some examples of application of the large-deviation approach to a variety of models are shortly discussed.
@@ 2 Simple and Biased Sampling
The most natural implementation of a stochastic system is to sample the states or configuration according to the original distribution of the system. Performing measurements means to take just the configurations, calculate the measurable quantities, and then obtain averages or histograms. This is introduced in the first subsection of this chapter. Next, as an example the Bernoulli process is considered. Then it is shown that one can reach the rare events, i.e., the tails of the distributions, by using biased sampling. Finally, this will be illustrated for the Bernoulli process again.
@@ 2.2 Simple Sampling for the Bernoulli Process
The result obtained within a simple simulation, here for $n=50$ and $\alpha=0.3$, could look like shown in Fig. 1. We consider three different values for the number $M$ of samples, namely $M=10^{4}$, $10^{6}$ and $10^{8}$. The histograms follow the true distribution Eq. (3) better for increasing $M$, but the tails are not accessible.
Figure 1: Numerically measured histograms of the number $l$ of 1’s obtained from $n=50$ coin flips. The histograms are measured from $M=10^{4}$, $10^{6}$ and $10^{8}$ experiments. The solid line shows the analytical solution Eq. (3).
@@ 2.4 Biased Sampling for the Bernoulli Process
In order to address these two problems, we first have to be able to sample according to more or less arbitrary distributions with possibly unknown normalization factor. This can be achieved by using Markov chains, which are covered next.
@@ 3 Markov Chain Monte Carlo Simulations
Often, the most simple numerical estimation procedure is uniform sampling. Here one generates a certain number $M$ states $\{\underline{y}^{i}\}$ ($i=1,\ldots,M$) randomly, with uniform probability. A typical value of $M$ could be, for example, $10^{6}$. For the Bernoulli example, one would obtain the uniform sampling by simply choosing for all entries $y_{i}=0,1$ with probability 0.5, instead of the actual probabilities $(1-\alpha)$ and $\alpha$, respectively. To estimate the average $\langle A\rangle$, one calculates the observable for all sampled states and weights the results with the correct probabilities. One has to normalize this weighted average by dividing by the sum of weights. This results in
$\langle A\rangle\approx\overline{A}^{(1)}:=\frac{\sum_{\underline{y}^{i}}A(\underline{y}^{i})P(\underline{y}^{i})}{\sum_{\underline{y}^{j}}P(\underline{y}^{j})}$ (12)
It can be easily checked that this way of normalizing is correct. Say the measurable quantity was boringly constant $A\equiv 1$, then Eq. (12) would result in $\sum_{\underline{y}^{i}}P(\underline{y}^{i})/$ $\left(\sum_{\underline{y}^{j}}P(\underline{y}^{j})\right)=1$ which is correct. Also, if the sample consisted of all $K$ possible states, then the sum $\sum_{\underline{y}^{i}}P(\underline{y}^{i})$ in the denominator is one by normalization and Eq. (12) reduces to the exact expectation value Eq. (11). Note also that for Eq. (12) it is sufficient to know the probabilities without normalization, since this cancels from the ratio.
@@
Note that $P$ denoted so far the original probabilities of the system. But, in the present context, it can be the biased probabilities $Q$ to drive the system to desired but unlikely configurations, as done for large-deviation sampling in section 2.3. In this case, one would have to chose $A$ appropriately to unbias the result, as discussed previously. We will come back to large-deviation sampling later in section 4, but for the moment we assume just arbitrary given probabilities $P$ according which we desire to sample, let them be the original or biased ones.
@@ 3.1 Markov Chains
Under specific conditions [4], as specified by the Theorems of Perron [5] and Frobenius [6], the probability distribution will converge to a stationary, i.e. time-independent distribution
@@
This will happen in particular if largest left eigenvalue $\lambda=1$ of the matrix $W$ has multiplicity 1, i.e., the corresponding eigenspace is one-dimensional.
The limit Eq. (16) has to be independent of any given initial state $\underline{y}(0)$. The Markov process is called ergodic in this case. This basically means that one can reach each state from any other state by a finite sequence of transitions. Note that there are non-ergodic systems, a simple example is shown in Fig. 4.
@@
In general and even if one takes the Metropolis-Hastings algorithm, there is a lot of freedom in choosing the transition probabilities, influencing strongly the efficiency of the corresponding algorithm. One general rule is, the more the configurations are changed between two time steps, the faster the algorithm will walk through configuration space, thus the more efficient the algorithm is. Unfortunately, these system-wide changes are hard to set up, often one has to be satisfied with single-variable changes.
@@ 3.2 MCMC for the Bernoulli Process
As mentioned before, there are in general many possibilities to set up the transition probabilities in the matrix $W$. Here we proceed as follows: Given the current state $\underline{y}=\underline{y}(t)$, we just select $n_{\rm c}$ times a randomly chosen entry and redraw it according to the original probabilities. Therefore, we perform $n_{\rm c}$ new coin tosses, while the outcomes of the other coin tosses contained in $\underline{y}$ are not touched. This is implemented by the following algorithm:
@@
The resulting configuration is $\underline{y}(t+1)$. Each time a new index $i$ is drawn randomly and independently, this might result in repeated redrawing some entries. Since each coin toss is independent, only the last one will be effective for a given entry. Thus, the actual number of entries actually redrawn may be smaller than $n_{\rm c}$. An implementation in C is contained [2] in mc_bernoulli.c [^4] The code can be compiled in a Unix shell by cc -o mc_bernoulli mc_bernoulli.c -lm . in the function bernoulli_mc_step0().
@@
For obtaining the transition probabilities, we consider two states $\underline{y}$ and $\underline{z}$ that are connected by a step in the Markov chain. Let $I_{\rm c}$ contain those indices where $\underline{y}$ and $\underline{z}$ differ, i.e., $I_{\rm c}=\{i|y_{i}\neq z_{i};i=1,\ldots,n\}$. Note that $\underline{y}$ and $\underline{z}$ can have a non-zero transition probability only when $n_{\rm c}\geq|I_{\rm c}|$
Furthermore, let $I_{\rm s}$ be the actual set of indices selected by the algorithm, with $|I_{\rm s}|\leq n_{\rm c}$. This set will most of the time be a super set of $I_{\rm c}$, because for some of the selected entries redoing the coin flip might lead to the previous outcome. Hence, for the entries in $I_{\rm s}\setminus I_{\rm c}$, the entries are not changed, like those entries which do not appear in $I_{\rm s}$ anyway. We denote by $R(I_{\rm s})$ the probability that the algorithm selects a certain subset $I_{\rm s}$, which might be a complicated function of $I_{\rm s}$. For the entries $i$ which appear in $I_{\rm s}$, the joint probability for obtaining the outcomes $z_{i}$ is simply $\prod_{i\in I_{\rm s}}\alpha^{z_{i}}(1-\alpha)^{1-z_{i}}$. Splitting $I_{\rm s}$ into $I_{\rm c}$ and $I_{\rm s}\setminus I_{\rm c}$, and summing over all possible sets $I_{\rm s}$, this leads to the following transition probability:
@@
Figure 6: Markov chain for the Bernoulli case with $n=50$ coin flips and $\alpha=0.5$. Shown is the number $l(t)=\sum_{i}y_{i}(t)$ of 1’s, averaged over 1000 runs, starting at $l(0)=0$. Three different numbers $n_{\rm c}=1,2,5,10$ of entry changes are considered. The horizontal line indicates the expectation value $\alpha n$.
In Fig. 6 $l(t)$ is shown for different number $n_{\rm c}$ of selected entries for changes. The more entries are changed, the faster the convergence is. This is in general true for all MCMC algorithms: the approach is the more efficient, the faster it can move through configuration space.
@@ 3.3 Metropolis-Hastings algorithm
Note that for the opposite move $\underline{z}\to\underline{y}$ the vectors $\underline{z}$ and $\underline{y}$ exchange and therefore $Q$ is replaced by $1/Q$. This also means that for one of the two moves between $\underline{z}$ and $\underline{y}$ the transition probability is one. One could use smaller acceptance probabilities, e.g. $1/2$ instead of 1 for one direction, the first term of the ${}\min\{\ldots\}$. But then one would have, to conserve detailed balance, to use the half of the probabilities for the opposite direction as well resulting in $\tilde{W}(\underline{y}\to\underline{z})=\min\Bigg{(}\frac{1}{2},\frac{1}{2}\frac{P(\underline{z})}{P(\underline{y})}\frac{A(\underline{z}\to\underline{y})}{A(\underline{y}\to\underline{z})}\Bigg{)}$, but this would reduce the overall number of performed changes. On the other hand, it is not possible to increase the acceptance rate 1 for the first term in the ${}\min\{\ldots\}$. In this sense, the Metropolis-Hastings choice is maximal with respect to the number of accepted moves, given the trial construction probabilities $A(\underline{z}\to\underline{y})$.
@@ 3.4 Metropolis-Hastings Algorithm for the Bernoulli Process
Hence, one can calculate the acceptance probability without performing the actual flip beforehand. Thus, one performs a single-variable flip with probability $\tilde{W}$ to yield $\underline{y}(t+1)$, otherwise the current configuration $\underline{y}(t)$ is kept. A corresponding C code [2] is contained in mc_bernoulli.c in the function bernoulli_metropolis().
@@ 4 Biased Markov Chain Sampling
Now we come back to the main target, to obtain for a random process the distribution $Q(S)$ of a measurable scalar quantity $S$, which we also call score. The score can be in principle any scalar function of the entire configuration $\underline{y}$. The distribution can be obtained in principle by considering all possible configurations $\underline{y}$ and adding up the probabilities $P({\underline{y}})$ of those configurations for which $S(\underline{y})$ has the desired value $S$, i.e.
@@ 4.1 Exponential Bias
where ${S}(\underline{y})$ is the score evaluated for the configuration. This bias appears very natural for physicists since it corresponds to the Boltzmann weight of the canonical ensemble. Therefore $\Theta$ is a temperature-like parameter, short just temperature, which controls the impact of the bias. Note that also other biases are possible, in principle everything that works is good. Nevertheless, for many systems the exponential bias works very well and it is also used for analytical calculations, where one can take advantage of biases to obtain tails of distributions in a similar way. For a short introduction, see Ref. [9]. Also more detailed texts, still well readable for physicists, are available [10, 11].
@@
  • In contrast to the physical canonical ensemble, also temperatures $\Theta<0$ make sense. Now, the bias is larger for configurations where $S=S(\underline{y})$ is larger than typical values of $S$.
@@
During the generation of configurations $\underline{y}$, one measures always the corresponding values $S=S(\underline{y})$, which finally can be collected in histograms which approximate the distributions $Q_{\Theta}(S)$ of $S$ in the biased ensemble at temperature $\Theta$. The results will depend of $\Theta$. This means that the histograms will cover a certain range of values of $S$. Therefore, one usually has to generate configurations for several values of $\Theta$ in order to obtain a full or at least a large range of values of $S$. Details on how the final estimate for the distribution $Q(S)$ can be obtained from combining the result generated at different values of $\Theta$ is presented in Sec. 4.3.
@@ 4.2 Bias for the Bernoulli Process
For the Bernoulli process of $n$ coin flips, the MCMC algorithm redraws in each step a number $n_{\rm c}$ of randomly chosen elements of the current configuration $\underline{y}(t)$, with the original coin probability $\alpha$ as of Eq. (1). For the changed elements, the current values are stored in a vector, i.e. an array old[]. This allows to restore the previous configuration, if the trial configuration is rejected. Note that for restoring the configuration, the current values are restored in reverse order, to prevent a mistake if some elements are by chance selected more than once for a change.
@@
algorithm biased-MC-Bernoulli($\underline{y}$, $n$, $n_{\rm mc}$, $n_{\rm c}$, $\Theta$, function $S(\cdot)$)
@@
Note that in case of acceptance, nothing has to be done, since the trial state is stored directly in $\underline{y}$. The C code [2] can be found in mc_bernoulli.c as function bernoulli_mc_step_bias().
@@
With the mc_bernoulli.c code [2] we can generate data for $S_{3+}$, The Metropolis algorithm is chosen by the -alg 2 option. First, we simulate the almost unbiased case for which we choose $\Theta=1000$. For this case, a large number $n_{\rm c}=45$ of changes are performed in each step. This is stated with the -change option when calling the program. Here, we consider $n=51$ coin flips for $\alpha=0.5$, which are the third and second last arguments, respectively. The number of Monte Carlo steps is the last argument, 100000 here. The first 10000 steps are for equilibration and therefore ignored for measuring the histogram, as stated by the argument for the -histo option. When calling the program in Unix from a command line, this can be achieved by entering
@@
You should perform simulations yourself using the code and observe how varying the parameters, like the number $n_{\rm c}$ of redrawn coin flips, lead to changes of the results. In particular the empirical acceptance probability is affected, which is also written into the output file. It is also interesting to change $\Theta$ and try to sample the largest possible values, which is here $S_{3+}=13$.